Using the definition of limit at infinity or infinite limits, prove that

a) lim𝑥 → 3 1/ (𝑥 − 3) 2 = ∞ 


1
Expert's answer
2022-03-19T02:39:39-0400

Let "C>0" be fixed. Solve

"\left| \frac{1}{\left( x-3 \right) ^2} \right|>C\Leftrightarrow \left| x-3 \right|<\frac{1}{\sqrt{C}}"

Thus for every C>0 there exists "\delta =\frac{1}{\sqrt{C}}" such that "\left| x-3 \right|<\delta \Rightarrow \left| \frac{1}{\left( x-3 \right) ^2} \right|>C"

By the definition

"\underset{x\rightarrow 3}{\lim}\frac{1}{\left( x-3 \right) ^2}=\infty"